Sol Summary – January 13th

Crew 290 Sol Summary Report 13-01-2024

Sol: 6

Summary Title: pHenomenal Day

Author’s name: Madelyn Hoying

Mission Status: Nominal

Sol Activity Summary:

The day kicked off with French toast and fried potatoes courtesy of Freebee (Anja) and Funk (Ben). The whole crew headed over to the Science Dome to prep EVA procedures and evaluate yesterday’s samples. There, PODO (Nicole) made the most exciting discovery of the day – a working pH meter, tucked away in the Science Dome! This is one that we can bring in the field as our experimental protocol had originally described, instead of requiring an in-hab workaround. We bumped our EVA back by an hour to give time to adjust the protocol and prepare the new sample setup. Chopper (Rebecca), Funk, Roots (Anna), and PODO set off on their EVA to Chandor Chasma with Melon (Madelyn) on CAPCOM and Freebee setting up the Science Dome. Two promising sample sites were collected, which the crew will continue analyzing in the coming days. Freebee is making a dinner of beans and rice, and the crew is looking forward to a rest day tomorrow.

Look Ahead Plan: Rest day, catch-up work time in the hab, team bonding

Anomalies in work: none

Weather: Windy, but less so than Sol 4

Crew Physical Status: nominal

EVA: 07 to Chandor Chasma, with Rebecca, Ben, Anna, and Nicole

Reports to be filed: Sol Summary, Journalist Report, Greenhab Report, Operations Report, EVA Report, Mid-Mission Report

Support Requested:

Additional CR2032 3V lithium batteries for the pH meter. The meter takes 4, and there are two meters in the Science Dome. If there are 8 batteries available, that would be great!
The temperature sensor broke in the field today, with a wire to the battery broken. Are there any 9V battery connectors available? If so, we can solder it to repair.

Copyright © The Mars Society. All rights reserved. | Main Site